OpenStudy (anonymous):

Find an equation of a circle that satisfies the given conditions. center (3,-1) radius 5

OpenStudy (amistre64):

(x-Cx)^2 + (y-Cy)^2=r^2

OpenStudy (amistre64):

\[(x-3)^2+(y+1)^2=25\]

OpenStudy (anonymous):

thank u

OpenStudy (amistre64):

yep